# Enhanced premelting at the ice–rubber interface using all-atom molecular dynamics simulation

### v1

@@ 3.3 Coupling of the dynamics of rubber molecules and premelting layers
4 Discussion
@@
Tribology of ice–rubber interface is influenced by multiple factors, including frictional heating, pressure melting, ice adhesion, and premelting layers, yet a comprehensive understanding of these effects remains elusive. Here, we aimed to clarify the rubber-induced effect on the ice premelting. From the atomistic molecular dynamics simulations, we have shown that the rubber, despite its hydrophobicity, induces structural disorder of the ice interface, enhancing the ice premelting. In contrast, based on the diffusion coefficient and liquid-like ratio, the premelting layer at the ice–rubber interface shows lower mobility compared to that at the ice–vapor interface. However, at temperatures approaching the melting point (267 to 269 K), a drastic increase in molecular dynamics was observed at the rubber-ice interface, accompanied by enhanced rubber molecular dynamics. These results demonstrate that ice and rubber mutually influence each other’s structure and dynamics, resulting in an ice-rubber interfacial layer with unique temperature-dependent properties.
@@
Why do hydrophobic polymers enhance premelting? Previous studies have proposed a depletion mechanism 47, 48, 26, in which water molecules are excluded from the hydrophobic polymer surface due to repulsive excluded volume effects, forming a water depletion layer. In contrast, our results show that water molecules can penetrate nanoscale voids within the polymer structures, where the roughness scale is comparable to the bilayer thickness of ice. This suggests that the nanoscale morphology of the polymer, via the excluded volume effect, disrupts hydrogen-bond networks of the water molecules within the premelting layers 49. Notably, the substantial increase in rubber flexibility observed in our study is likely triggered by bilayer melting, which occurs a few kelvins below the melting point at the ice-vapor interface 14. As a result, the topmost layer becomes a mixed rubber–ice state, encouraging lower layers (Layer 2 and Layer 3) to behave as if they are the topmost interface layer.
@@
Ice–rubber tribology is a multiscale phenomenon, and our work provides fundamental molecular-level insights into the rubber–ice interface. We speculate that the premelting layers influence both the effective friction coefficient and adhesion to ice. Bridging these molecular-level understanding with macroscopic observations may contribute to the development of functional polymer materials and their industrial applications.
@@ Acknowledgments
Figure S 1: Density profiles of water oxygen atom (Ice O) and rubber atoms (Rubber) at 254, 259, 264 and 269 K.
Temperature [K] Vapor Rubber
254 0.854 0.771
259 0.862 0.810
264 0.861 0.792
267 0.878 0.865
269 0.892 0.876
Table S 1: Diffusion exponent, $\alpha$ ($\rm MSD\propto t^{\alpha}$), at ice–vapor interface (Vapor) and ice–rubber interface (Rubber).

### v2

@@ 3.4 Coupling of the dynamics of rubber molecules and premelting layers
4 Conclusions
Tribology of ice–rubber interface is influenced by multiple factors, including frictional heating, pressure melting, ice adhesion, and premelting layers, yet a comprehensive understanding of these effects remains elusive at atomistic resolution. In previous studies on ice or liquid–water interfaces, idealized, atomically smooth hydrophobic walls were employed, revealing depletion effects and enhanced water mobility janecek2007interfacial, baran2022ice. In contrast, recent studies have modeled ice interfaces using fully atomistic surface representations of materials such as graphene and polymers, showing that atomistically heterogeneous surfaces enable site-specific interactions between water and particular surface atoms janecek2007interfacial, sato2025molecular, but such effects cannot be captured by smooth-wall models. However, no prior molecular simulation has examined ice in contact with an atomistic rubber surface models, leaving the molecular origins of rubber-induced premelting entirely unexplored.
@@ 4 Conclusions
In this study, we employed an atomistic rubber model for the interface, demonstrating that rubber, despite its hydrophobicity, induces structural disorder at the ice interface, thereby enhancing ice premelting. In contrast, based on the diffusion coefficient and liquid-like ratio, the premelting layer at the ice–rubber interface shows lower mobility compared to that at the ice–vapor interface. Notably, at temperatures approaching the melting point (267 to 269 K), a drastic increase in molecular dynamics was observed at the rubber-ice interface, accompanied by enhanced rubber molecular dynamics. These results demonstrate that ice and rubber mutually influence each other’s structure and dynamics, resulting in an ice-rubber interfacial layer with unique temperature-dependent properties.
@@
Why do hydrophobic polymers enhance premelting? Previous studies have proposed a depletion mechanism janecek2007interfacial, poynor2006water, chen2017icephobic, in which water molecules are excluded from the hydrophobic polymer surface due to repulsive excluded volume effects, forming a water depletion layer. In contrast, our results show that water molecules can penetrate nanoscale voids within the polymer structures (Fig. 4a), where the thickness of roughness is comparable to the bilayer thickness of ice. This suggests that the nanoscale morphology of the polymer, via the excluded volume effect, disrupts hydrogen-bond networks of the water molecules within the premelting layers (Fig. S3kimmel2009no. Notably, the substantial increase in rubber flexibility observed in our study is likely triggered by bilayer melting, which occurs a few kelvins below the melting point at the ice-vapor interface sanchez2017experimental. As a result, the topmost layer becomes a mixed rubber–ice state, encouraging lower layers (Layer 2 and Layer 3) to behave as if they are the topmost interface layer.
Our observations of molecular structure and dynamics can be connected to experimental results. Surface force apparatus measurements have shown a sharp decrease in viscosity of the premelting layers above $5~\mathrm{{}^{\circ}C}$ lecadre2020ice. Because the viscosity and diffusivity are inversely related thorough the Stokes-Einstein equation baran2022ice, these experimental measurements correspond to the increase in the diffusion coefficient of the premelting layers in our simulations. In contrast, the apparent contact area of the hydrophobic polymer with ice changed little in experiment pallbo2024enhanced, although we observed increased structural disordering. This suggests that the increased disordering of premelting layers at the rubber interface observed in our simulations are confined to only a few molecular layers, making them detectable at the molecular scale but likely averaged out at the spatial resolution of the experimental measurements.
Our simulation setup employed a minimal model, consisting of a single SBR rubber composition in contact with the basal face of ice. In practical applications, however, rubber formulations typically vary in composition and contain cross-links, while multiple ice crystal faces may be exposed depending on environmental conditions. Our study serves as a baseline system that clarifies key physical phenomena, such as confined premelting dynamics and the coupled interfacial dynamics between ice and rubber chains. This framework can be systematically extended in future studies to incorporate additional material complexities, including fillers, copolymer architectures, surface treatments, and different ice crystal faces.
@@
Ice–rubber tribology is a multiscale phenomenon, and our work provides fundamental molecular-level insights into the rubber–ice interface. Although we have focused on the equilibrium properties, in real-life conditions, the ice interface is often subjected to sliding, and under such conditions, the effective viscosity of the liquid decreases due to shear thinning zhao2022new, baran2022ice, baran2024confinement. We speculate that the premelting layers influence both the effective friction coefficient and adhesion to ice. Bridging these molecular-level understanding with macroscopic observations may contribute to the development of functional polymer materials and their industrial applications.
Supporting Information
Method for calculating orientational order parameters; a table of diffusion coefficients and scaling exponents; additional figures including a snapshot of the full simulation box, density profiles, orientational order parameters, mean square displacements, diffusion coefficients, $g(x)$ distributions, and liquid-like ratio.
@@ Acknowledgments
4.1 Orientation order parameter
To quantify the structural ordering of the ice premelting layer, we computed the orientational order parameter and the intensity of the peak sanchez2017experimental. The orientational order parameter was defined as,
$\cos{\theta}=\frac{\bm{n}_{z}\cdot\bm{r}_{\rm{OH}}}{|\bm{r}_{\rm{OH}}|}$ (5)
where $\bm{n}_{z}$, $\bm{r}_{\rm{OH}}$ are the normalized vector in the surface direction and the vector from the oxygen atom to the dummy charge.
Temperature Vapor Rubber
[K] $\alpha$ Diffusion Coefficient $\mathrm{m^{2}/s}$] $\alpha$ Diffusion Coefficient $\mathrm{m^{2}/s}$]
254 0.922 $1.817\times 10^{-10}$ 0.778 $4.548\times 10^{-11}$
259 0.949 $2.649\times 10^{-10}$ 0.808 $6.704\times 10^{-11}$
264 0.953 $3.596\times 10^{-10}$ 0.825 $9.828\times 10^{-11}$
267 0.958 $4.187\times 10^{-10}$ 0.860 $1.349\times 10^{-10}$
269 0.963 $4.480\times 10^{-10}$ 0.854 $1.606\times 10^{-10}$
Table S1: Diffusion exponent (2D, Layer1), $\alpha$ ($\rm MSD\propto t^{\alpha}$), and diffusion coefficient (2D, Layer1) at the ice–vapor interface (Vapor) and the ice–rubber interface (Rubber).
Figure S1: Snapshot showing the full simulation box.
Figure S2: Density profiles of water oxygen atom (Ice O) and rubber atoms (Rubber) at 254, 259, 264, 267 and 269 K.
Figure S3: Probability density function of the orientational order parameter, $P(\theta)$, defined as the angle, $\theta$, between the dipole moment and the interface normal direction.
Figure S4: Peak values of the probability density function for the orientational order parameter at $\cos\theta\approx 0.31$, where the peak position was determined from Layer 4 at each temperature.
Figure S5: Parallel diffusion coefficient calculated for molecules located in Layer 1 (Layer 1) and in the combined Layers 1 and 2 (Layer 1–2) at both ice–vapor and ice–rubber interfaces.
Figure S6: Density profiles at the ice–rubber interface under load pressures of $0.1~\mathrm{MPa}$ and $1~\mathrm{MPa}$ at 254 and 269 K.
Figure S7: Diffusive molecular dynamics at $0.1~\mathrm{MPa}$ in comparison with $1~\mathrm{MPa}$. (a) MSD of parallel diffusion. (b) Parallel diffusion coefficient. Simulations at the rubber–ice interface under a load pressure $1~\mathrm{MPa}$ yielded ice–vapor ($0.1~\mathrm{MPa}$) and ice–rubber ($1~\mathrm{MPa}$), while that under a load pressure of $0.1~\mathrm{MPa}$ yielded ice–rubber ($0.1~\mathrm{MPa}$).
Figure S8: Profiles of $g(x)$ for ice–rubber interfaces at 254, 259, 264, and 267 K. Dashed black lines indicate the threshold used to distinguish solid-like and liquid-like molecules.
Figure S9: Liquid ratio in Layer 4 based on $g(x)$. Error bars show the standard error of the mean.
{tocentry}[Uncaptioned image]